OpenStudy (prince1342):

WILL FAN AND MEDAL Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube?

OpenStudy (prince1342):

This confuses me because I dont know if I should use the formula for a square or a pyramid or cube

OpenStudy (prince1342):

@mathstudent55 Can you please help me?

OpenStudy (prince1342):

@welshfella

OpenStudy (welshfella):

compare the formulae for the volume of the cube and the volume of a pyramid

OpenStudy (prince1342):

Hmm ok

OpenStudy (prince1342):

I got 12 cm

OpenStudy (welshfella):

Hint write the volume of a cube as s^2 * h where s= length of 1 side = h

OpenStudy (mathmale):

I'd suggest you sketch a cube and a pyramid separately, and then draw the pyramid inside the cube. If the volume of the cube is 36 cm^3, what is the length of one side of the cube? What is the height of the pyramid? What is the area of the base of the pyramid?

OpenStudy (welshfella):

in the case of a cube the length of ecah side is equal to its height

OpenStudy (prince1342):

Ohhh so would 12 cm be it then?

OpenStudy (welshfella):

yes because Area cube = s^2 h and area of pyramis = 1/3 s^2 h

OpenStudy (prince1342):

And what does the s stand for?

OpenStudy (welshfella):

the side of cube which will equal the side of the pyramid's base

OpenStudy (prince1342):

ohhh okay

OpenStudy (prince1342):

I have another question that I need help on

OpenStudy (welshfella):

- becaus eit fits exactly into the cube.

OpenStudy (welshfella):

plz post it separately

OpenStudy (prince1342):

Yeah I'm going to, that way I can give you another medal

OpenStudy (welshfella):

ok

OpenStudy (mathmale):

welshfella: "because Area cube = s^2 h" Please think about this and make the necessary correction. How would you describe a "cube"?

OpenStudy (welshfella):

yes I should have said volume